# Detector noise in continuous-variable quantum key distribution

### v2

@@ Abstract
Detector noise is a critical factor in determining the performance of a quantum key distribution (QKD) system. In continuous-variable (CV) QKD with optical coherent detection, the widely adopted trusted detector noise model improves both the secret key rate and transmission distance. This model assumes that detector noise is inherently random and inaccessible to an adversary. Its validity depends on two key assumptions: (1) the detector can be accurately calibrated by the legitimate user and remains isolated from the adversary, and (2) the detector noise is truly random from the adversary’s perspective. While substantial research has focused on strengthening the first assumption, the second is far more difficult to justify. Indeed, in quantum random number generation using optical homodyne detection, detector noise is often treated as “untrusted” and unsuitable for generating true randomness. In this paper, we introduce a calibrated detector noise model for CV-QKD, which relies solely on the first assumption. Specifically, our model applies even when detector noise is predictable to the adversary. We analyze the electrical noise of a commercial balanced photoreceiver and perform numerical simulations to compare different noise models. Our results show that when the detector noise variance is an order of magnitude below the vacuum noise, the proposed model achieves a secret key rate comparable to that of the trusted detector noise model, while eliminating the questionable assumption of “truly random” detector noise.
@@ I Introduction
One important application of quantum information science is quantum key distribution (QKD), a protocol that enables two remote parties (Alice and Bob) to establish cryptographic keys with proven security [1, 2, 3, 4].
@@
In practice, no QKD implementation is perfect. Real-world imperfections introduce system noise even in the absence of Eve. If the legitimate users cannot distinguish intrinsic system noise from noise caused by Eve, a conservative approach is to attribute all observed noise to Eve’s attack and estimate her potential information gain accordingly. While this “untrusted noise” model may overestimate Eve’s information and reduce QKD performance, it minimizes reliance on additional assumptions about the system. Consequently, it is commonly used in discrete-variable (DV) QKD based on single-photon detection [5], where the system noise is at a low level.
@@
In principle, the untrusted noise model described above can also be applied to continuous-variable (CV) QKD based on coherent detection, a potentially cost-effective solution compatible with classical fiber-optic networks [6, 7, 8]. However, this approach often leads to poor performance, particularly under high channel loss. The primary limiting factor is the electrical noise of the optical coherent detector, which is typically an order of magnitude lower than vacuum noise [9, 10, 11]. In such cases, attributing detector noise to Eve’s attack significantly reduces both the secret key rate and the achievable QKD distance.
@@
One approach to mitigate this issue is to adopt the trusted detector noise model, which is widely used in practical CV-QKD implementations [8, 12, 13, 14, 15, 16]. In this model, detector noise is treated as an intrinsic property that cannot be controlled or accessed by Eve. By distinguishing such intrinsic detector noise from other untrusted noise—potentially introduced by Eve—and treating them differently in the security analysis, a tighter bound on information leakage can be obtained, leading to improved QKD performance. The trade-off, however, is that the legitimate users must make additional assumptions about the QKD system.
@@
The validity of the trusted detector noise model relies on two key assumptions: (1) the detector can be accurately calibrated by the legitimate user and remains isolated from the adversary, and (2) the detector noise is truly random from the adversary’s perspective. While substantial research has focused on strengthening the first assumption [17, 18, 19, 20], the second is far more difficult to justify. In principle, only quantum noise can be considered truly random and thus suitable for cryptographic applications. However, verifying the quantum origin of the electrical noise in a real-life optical coherent detector remains challenging. In fact, in quantum random number generation (QRNG) using optical homodyne detection, detector noise is often treated as “untrusted” and unsuitable for generating true randomness [21, 22, 23, 24, 25].
@@
In this paper, we introduce a calibrated detector noise model for CV-QKD, which relies solely on the first assumption. Specifically, our model applies even when detector noise is predictable to the adversary. We analyze the electrical noise of a commercial balanced photoreceiver and perform numerical simulations to estimate the asymptotic secret key rates of the Gaussian-modulated coherent states (GMCS) QKD protocol [8] under different noise models. Our results show that the proposed model can achieve a secret key rate comparable to that of the trusted-noise model, while removing the questionable assumption of “truly random” detector noise.
@@ II Trusted and untrusted detector noise models in CV-QKD
The basic procedures of the GMCS QKD without switching [26] are as follows: for each transmission, Alice draws two random numbers $x_{A}$ and $p_{A}$ and transmits the coherent state $|x_{A}+ip_{A}\rangle$ to Bob through an insecure quantum channel. Here $x_{A}$ and $p_{A}$ are Gaussian random numbers with zero mean and a variance of $V_{A}N_{0}$, where $V_{A}$ is the modulation variance and $N_{0}$ = 1/4 denotes the shot-noise variance. Upon receiving the state, Bob performs a simultaneous measurement of both quadratures X and P using a conjugate homodyne detection setup. By repeating this quantum communication process multiple times, Alice and Bob collect sufficient data to estimate the channel transmittance and the level of excess noise via an authenticated classical channel. If the observed noise is below a predetermined security threshold, they proceed with key extraction by applying error reconciliation followed by privacy amplification, ultimately generating a shared secret key.
@@
Given the above parameters, the secret key rate of the GMCS QKD protocol can be determined. In this paper, we consider the protocol using reverse reconciliation and conjugate homodyne detection scheme. The general formula for the lower bound of the asymptotic key rate is then given by [27, 12]
@@
In general, calculating the Holevo bound $\chi_{BE}$ is nontrivial and depends on the specific noise model adopted. Under collective attacks, the Holevo bound for GMCS QKD is given by [12, 28]
@@
where $G(x)=(x+1)\log_{2}(x+1)-x\log_{2}x$. The values of $\lambda_{i}$ are given by [12, 28]
@@
Note in Eq. 6 the channel noise and the detection noise depend on the noise model employed. In the trusted detector noise model, Eve has no control over Bob’s setup, and the noises of Bob’s detectors are assumed to be truly random to Eve. The channel noise $\Xi_{\text{ch}}^{\text{trusted}}$ therefore only depends on the channel transmittance $T$ and noises outside Bob’s detector, and is given by [29]
@@
where $\xi_{A}$ represents both the excess noise contributed by Alice during her state preparation and the excess noise from the channel. Similarly, the detection noise $\Xi_{\text{det}}^{\text{trusted}}$ on Bob’s side relies on the detector efficiency and the electrical noise of the detector. Assuming conjugate homodyne detection scheme, the detection noise is given by [28]
@@ III Experimental characterization of optical homodyne detector
The differential output of the balanced photoreceiver represents the quadrature value of the vacuum state, from which perfect randomness could be generated [21]. Shown in FIG.2 is the total variance of the detector output, $\sigma_{T}^{2}$, that was measured at different LO powers, for different detector setups (bandwidth of 20 MHz or 500 MHz, gain of $5\times 10^{3}$ V/A or $10\times 10^{3}$ V/A).
@@
Assuming that the vacuum noise is independent of the electrical noise of the detector [30], the relationship among the three noise variances can be expressed as:
@@
In QRNGs using optical homodyne detection, the vacuum noise is typically treated as quantum noise, while the detector’s electrical noise is assumed to be classical noise. To quantify their relative contributions, the quantum-to-classical noise ratio (QCNR) is defined as [30]:
@@ V Numerical simulation of secret key rates
We perform numerical simulations of the key rates for the GMCS QKD protocol under the three detector noise models, with the reconciliation efficiency $f=0.95$ and the detector efficiency $\eta_{B}=0.5$. We plot the key rates over 150km with fiber attenuation 0.20.2dB/km. For each fiber length, we optimize the key rate over the modulation variance $V_{A}$. To account for phase noise in a coherent communication system, the excess noise is modeled as $\xi_{A}=0.01+0.01V_{A}$ [31]. FIG. 5 plots the key rates for different electrical noises of the detector: $\nu_{B}=0.01,0.1,0.28$.
Figure 5: Key rates for GMCS QKD. The three different detector noise models are simulated: trusted, untrusted, and calibrated. The key rates are optimized over the modulation variance $V_{A}$. Other simulation parameters are: $f=0.95$, $\eta_{B}=0.5$, and $\nu_{B}=0.01,0.1,0.28$. The fiber attenuation used is 0.20.2dB/km.
The case of $\nu_{B}=0.01$ corresponds to a homodyne detector with very low noise. Achieving such low noise typically requires a detector with a narrow bandwidth [32]. As shown in FIG. 5, although our calibrated detector noise model is based on more realistic assumptions than the trusted detector noise model, the resulting loss in key rate is minimal. Moreover, both the trusted and calibrated models yield significantly higher key rates compared to the untrusted noise model. The case of $\nu_{B}=0.1$ represents more realistic detector noise in high-speed CV-QKD systems [9, 10, 11] while $\nu_{B}=0.28$ corresponds to the detector described in Section III when operated at 500MHz ($\text{QCNR}=5.5$). In these regimes, the key rates for the untrusted detector noise model completely vanish, and the difference between our calibrated noise model and the trusted noise model becomes more visible. $\nu_{B}=0.1$ yields similar key rates at short distances and deviates slightly at long distances, while $\nu_{B}=0.28$ shows a significant drop in key rates after 100 km. These simulation results indicate that the calibrated noise model achieves a well-balanced trade-off between key rate and security.
@@ VI Discussion
In this paper, we introduce a third option—the calibrated detector noise model—which lies between the untrusted and trusted noise models. This model relies solely on the first assumption: that the detector is isolated from Eve’s control. It allows for the possibility that detector noise is fully predictable to Eve. A key advantage is that it eliminates the need to validate the true randomness of detector noise, a task that can be challenging in practice. The trade-off is a slight reduction in QKD performance compared to the trusted noise model. Nevertheless, our numerical simulations show that the resulting loss in secret key rate is minimal under practical parameters. Moreover, the new model can be readily incorporated into existing CV-QKD experimental setups.
Acknowledgments We thank an anonymous reviewer for pointing out a mistake in the autocorrelation calculation in an earlier version of the manuscript. Bing Qi acknowledges support from NYU-Shanghai start-up funds.

### v3

@@ Abstract
In continuous-variable (CV) QKD with optical coherent detection, the widely adopted trusted detector noise model improves both the secret key rate and the transmission distance. This model assumes that detector noise is inherently random and inaccessible to an adversary. While substantial research has focused on shielding the detector, it is far more difficult to justify the adversary’s ignorance of the detector noise. In this paper, we introduce a calibrated detector noise model for CV-QKD, which relies solely on the isolation of the detector from the adversary’s intervention. Specifically, our model applies even when detector noise is predictable to the adversary. We analyze the electrical noise of a commercial balanced photoreceiver and perform numerical simulations to compare different noise models. Our results show that when the detector noise variance is an order of magnitude below the vacuum noise, the proposed model achieves a secret key rate comparable to that of the trusted detector noise model, while eliminating the questionable assumption of “truly random” detector noise.
@@ I Introduction
One important application of quantum information science is quantum key distribution (QKD), a protocol that enables two remote parties (Alice and Bob) to establish cryptographic keys with proven security [gisin2002quantum, diamanti2016practical, xu2020secure, pirandola2020advances].
@@
In practice, no QKD implementation is perfect. Real-world imperfections introduce system noise even in the absence of Eve. If the legitimate users cannot distinguish intrinsic system noise from noise caused by Eve, a conservative approach is to attribute all observed noise to Eve’s attack and estimate her potential information gain accordingly. While this “untrusted noise” model may overestimate Eve’s information and reduce QKD performance, it minimizes reliance on additional assumptions about the system. Consequently, it is commonly used in discrete-variable (DV) QKD based on single-photon detection [bennett2014quantum], where the system noise is at a low level.
@@
In principle, the untrusted noise model described above can also be applied to continuous-variable (CV) QKD based on coherent detection, a potentially cost-effective solution compatible with classical fiber-optic networks [ralph1999continuous, hillery2000quantum, grosshans2003quantum]. However, this approach often leads to poor performance, particularly under high channel loss. The primary limiting factor is the electrical noise of the optical coherent detector, which is typically an order of magnitude lower than vacuum noise [Chi2011, huang2013300, kumar2012versatile]. In such cases, attributing detector noise to Eve’s attack significantly reduces both the secret key rate and the achievable QKD distance.
@@
One approach to mitigate this issue is to adopt the trusted detector noise model, which is widely used in practical CV-QKD implementations [grosshans2003quantum, lodewyck2007quantum, qi2007experimental, jouguet2013experimental, huang2016long, zhang2019continuous, hajomer2025chipbasedcvqkd, wang2025high]. In this model, detector noise is treated as an intrinsic property that cannot be controlled or accessed by Eve. By distinguishing such intrinsic detector noise from other untrusted noise—potentially introduced by Eve—and treating them differently in the security analysis, a tighter bound on information leakage can be obtained, leading to improved QKD performance. The trade-off, however, is that the legitimate users must make additional assumptions about the QKD system.
@@
The validity of the trusted detector noise model relies on two key assumptions: (1) the detector can be accurately calibrated by the legitimate user and remains isolated from the adversary, and (2) the detector noise is truly random from the adversary’s perspective. While substantial research has focused on strengthening the first assumption [jouguet2013preventing, zhang2020one, brunner2020precise, zou2022rigorous], the second is far more difficult to justify. In principle, only quantum noise can be considered truly random and thus suitable for cryptographic applications. However, verifying the quantum origin of the electrical noise in a real-life optical coherent detector remains challenging. In fact, in quantum random number generation (QRNG) using optical homodyne detection, detector noise is often treated as “untrusted” and unsuitable for generating true randomness [gabriel2010generator, qi2010high, ma2013postprocessing, zheng20196, bai202118, zhang2024selftestingqrng, wang2025highly].
@@
In this paper, we introduce a calibrated detector noise model for CV-QKD, which relies solely on the first assumption. Specifically, our model applies even when detector noise is predictable to the adversary. We analyze the electrical noise of a commercial balanced photoreceiver and perform numerical simulations to estimate the asymptotic secret key rates of the Gaussian-modulated coherent states (GMCS) QKD protocol [grosshans2003quantum] under different noise models. Our results show that the proposed model can achieve a secret key rate comparable to that of the trusted-noise model, while removing the questionable assumption of “truly random” detector noise.
@@ II Trusted and untrusted detector noise models in CV-QKD
The basic procedures of the GMCS QKD without switching [weedbrook2004quantum] are as follows: for each transmission, Alice draws two random numbers $x_{A}$ and $p_{A}$ and transmits the coherent state $|x_{A}+ip_{A}\rangle$ to Bob through an insecure quantum channel. Here $x_{A}$ and $p_{A}$ are Gaussian random numbers with zero mean and a variance of $V_{A}N_{0}$, where $V_{A}$ is the modulation variance and $N_{0}$ = 1/4 denotes the shot-noise variance. Upon receiving the state, Bob performs a simultaneous measurement of both quadratures X and P using a conjugate homodyne detection setup. By repeating this quantum communication process multiple times, Alice and Bob collect sufficient data to estimate the channel transmittance and the level of excess noise via an authenticated classical channel. If the observed noise is below a predetermined security threshold, they proceed with key extraction by applying error reconciliation followed by privacy amplification, ultimately generating a shared secret key.
@@
Given the above parameters, the secret key rate of the GMCS QKD protocol can be determined. In this paper, we consider the protocol using reverse reconciliation and conjugate homodyne detection scheme. The general formula for the lower bound of the asymptotic key rate is then given by [Devetak2005, lodewyck2007quantum]
@@
In general, calculating the Holevo bound $\chi_{BE}$ is nontrivial and depends on the specific noise model adopted. Under collective attacks, the Holevo bound for GMCS QKD is given by [lodewyck2007quantum, Fossier2009]
@@
where $G(x)=(x+1)\log_{2}(x+1)-x\log_{2}x$. The values of $\lambda_{i}$ are given by [lodewyck2007quantum, Fossier2009]
@@
Note in Eq. 6 the channel noise and the detection noise depend on the noise model employed. In the trusted detector noise model, Eve has no control over Bob’s setup, and the noises of Bob’s detectors are assumed to be truly random to Eve. The channel noise $\Xi_{\text{ch}}^{\text{trusted}}$ therefore only depends on the channel transmittance $T$ and noises outside Bob’s detector, and is given by [Scarani2009]
@@
where $\xi_{A}$ represents both the excess noise contributed by Alice during her state preparation and the excess noise from the channel. Similarly, the detection noise $\Xi_{\text{det}}^{\text{trusted}}$ on Bob’s side relies on the detector efficiency and the electrical noise of the detector. Assuming conjugate homodyne detection scheme, the detection noise is given by [Fossier2009]
@@ III Experimental characterization of optical homodyne detector
The differential output of the balanced photoreceiver represents the quadrature value of the vacuum state, from which perfect randomness could be generated [gabriel2010generator]. Shown in FIG.2 is the total variance of the detector output, $\sigma_{T}^{2}$, that was measured at different LO powers, for different detector setups (bandwidth of 20 MHz or 500 MHz, gain of $5\times 10^{3}$ V/A or $10\times 10^{3}$ V/A).
@@
Assuming that the vacuum noise is independent of the electrical noise of the detector [Feihu2012], the relationship among the three noise variances can be expressed as:
@@
In QRNGs using optical homodyne detection, the vacuum noise is typically treated as quantum noise, while the detector’s electrical noise is assumed to be classical noise. To quantify their relative contributions, the quantum-to-classical noise ratio (QCNR) is defined as [Feihu2012]:
@@
It is also worth noting that our experiment presents the basic operation of Bob’s detector calibration process. In practice, to deal with unstable detector noises, the calibration could be done in real time by randomly blocking a chunk of signals during the QKD process.
@@ V Numerical simulation of secret key rates
We perform numerical simulations of the key rates for the GMCS QKD protocol under the three detector noise models, with the reconciliation efficiency $f=0.95$ and the detector efficiency $\eta_{B}=0.5$. We plot the key rates over 150km with fiber attenuation 0.20.2dB/km. For each fiber length, we optimize the key rate over the modulation variance $V_{A}$. To account for phase noise in a coherent communication system, the excess noise is modeled as $\xi_{A}=0.01+0.01V_{A}$ [qi2015generating]. FIG. 5 plots the key rates for different electrical noises of the detector: $\nu_{B}=0.01,0.1,0.28$.
Figure 5: Key rates for GMCS QKD. The three different detector noise models are simulated: trusted (dashed line), untrusted (dash-dotted line), and calibrated (solid line). The electrical noises of the detector are $\nu_{B}=0.01$ (blue), $\nu_{B}=0.1$ (red), and $\nu_{B}=0.28$ (black). The zoom-in subplot shows the key rates in the low noise regime $\nu_{B}=0.01,0.1$, within the typical length of access networks (25km-50km). The key rates are optimized over the modulation variance $V_{A}$. Other simulation parameters are: $f=0.95$, $\eta_{B}=0.5$. The fiber attenuation used is 0.20.2dB/km.
The case of $\nu_{B}=0.01$ corresponds to a homodyne detector with very low noise. Achieving such low noise typically requires a detector with a narrow bandwidth [tian2022experimental]. As shown in FIG. 5, although our calibrated detector noise model is based on more realistic assumptions than the trusted detector noise model, the resulting loss in key rate is minimal. Moreover, both the trusted and calibrated models yield significantly higher key rates compared to the untrusted noise model. The case of $\nu_{B}=0.1$ represents more realistic detector noise in high-speed CV-QKD systems [Chi2011, huang2013300, kumar2012versatile] while $\nu_{B}=0.28$ corresponds to the detector described in Section III when operated at 500MHz ($\text{QCNR}=5.5$). In these regimes, the key rates for the untrusted detector noise model completely vanish, and the difference between our calibrated noise model and the trusted noise model becomes more visible. $\nu_{B}=0.1$ yields similar key rates at short distances and deviates slightly at long distances, while $\nu_{B}=0.28$ shows a significant drop in key rates after 100 km. These simulation results indicate that the calibrated noise model achieves a well-balanced trade-off between key rate and security.
@@ VI Discussion
In this paper, we introduce a third option—the calibrated detector noise model—which lies between the untrusted and trusted noise models. This model relies solely on the first assumption: that the detector is isolated from Eve’s control. It allows for the possibility that detector noise is fully predictable to Eve. A key advantage is that it eliminates the need to validate the true randomness of detector noise, a task that can be challenging in practice. The trade-off is a slight reduction in QKD performance compared to the trusted noise model. Nevertheless, our numerical simulations show that the resulting loss in secret key rate is minimal under practical parameters. Specifically, when the detector noise variance is an order of magnitude below the vacuum noise ($\nu_{B}=0.1$), the proposed model achieves a secret key rate comparable to that of the trusted detector noise model. We also simulate the performance of a characterized commercial balanced detector ($\nu_{B}=0.28$), and show that even at a relatively high detector noise level, the calibrated model can still tolerate up to 20 dB of channel loss. Moreover, the new model can be readily incorporated into existing CV-QKD experimental setups.
Acknowledgments We thank an anonymous reviewer for pointing out a mistake in the autocorrelation calculation in an earlier version of the manuscript. We thank Shlok Nahar for comments on the trusted and untrusted detector noise models. Bing Qi acknowledges support from NYU-Shanghai start-up funds.